OpenStudy (anonymous):

solve a-2b=2 a+b=2(b+3)

OpenStudy (anonymous):

@BVB4rmy

OpenStudy (anonymous):

a-2b=2 ----(1) a+b=2(b+3) ----(2) from (1) a = 2 + 2b put in (2) a+b=2(b+3) 2+2b+b=2(b+3) 2+3b =2b+6 3b-2b =6-2 b = 4 put in (1) a-2b=2 a-2(4)=2 a-8=2 a=2+8 a=10 S.S= {a,b}={10,4}